OpenStudy (anonymous):

A kite rises at an angle of 14° with the ground. Find the distance the kite has flown (hint: the length of the string) when it has covered a horizontal distance of 150 feet. Round your answer to the nearest foot.

OpenStudy (anonymous):

150/Cos 14

OpenStudy (anonymous):

cos(14°) = 150/x x = 150/cos(14°) = 154.6 x = 155 ft